40. The demand equation for a product is given by P = 30 – 0.1Q2
i) Write an equation for the point elasticity as a function of quantity.
ii) At what price is demand unitary elastic?
1
Expert's answer
2020-11-10T10:03:51-0500

i)"Qp=\sqrt{\frac{P-30}{0.1}}"

"Ep=Q'p\frac{P}{Q}"


"Ep=(\sqrt{\frac{P-30}{0.1}})'\frac{P}{\sqrt{\frac{P-30}{0.1}}}=(2\sqrt{\frac{P-30}{0.1}})^{-1}(\frac{P}{\sqrt{\frac{P-30}{0.1}}})=\frac{P}{2P-30}"


"Ep="

ii) Ep=1


"1=\frac{P}{2P-30}"

P=30